OpenStudy (help!!!!):

A fraction equals 8/9. The numerator is 2 more than 6 times the quotient of the denominator and 7. Find the reciprocal of the original fraction

OpenStudy (help!!!!):

@jim_thompson5910

OpenStudy (help!!!!):

So this is my set up. 6(x+7)=8/9 or is it 6(x+7)/x=8/9?

jimthompson5910 (jim_thompson5910):

`A fraction equals 8/9` so we have 2 numbers x and y divide like this x/y and they will reduce to 8/9 so that means x/y = 8/9 agreed?

OpenStudy (help!!!!):

Yup but could we approach this with 1 variable?

jimthompson5910 (jim_thompson5910):

yes, but it helps to start off with 2 variables then do a substitution

jimthompson5910 (jim_thompson5910):

x = numerator y = denominator `The numerator is 2 more than 6 times the quotient of the denominator and 7` so (numerator) = 6*(denominator/7) + 2 x = 6*(y/7) + 2 agreed?

OpenStudy (help!!!!):

agree!

jimthompson5910 (jim_thompson5910):

so we go from \[\Large \frac{x}{y} = \frac{8}{9}\] to \[\Large \frac{6(y/7)+2}{y} = \frac{8}{9}\]

OpenStudy (help!!!!):

Alrighty! Thank you

OpenStudy (help!!!!):

I got confused with the wording of the equation :)

jimthompson5910 (jim_thompson5910):

you're welcome